You have just been hired by FAB Corporation, the manufacturer of a revolutionary new garage door opening device. The president has asked that you review the company's costing system and "do what you can to help us get better control of our manufacturing overhead costs." You find that the company has never used a flexible budget, and you suggest that preparing such a budget would be an excellent first step in overhead planning and control. After much effort and analysis, you determined the following cost formulas and gathered the following actual cost data for March Actual Cost in Cost Formula March Utilities $16,400 plus $0.11 per machine-hour $38,900 plus $2.00 per machine-hour $0.50 per machine-hour 20,400 $72,700 $ 9,800 $120,900 $ 70,000 Maintenance Supplies Indirect labor $94,000 plus $1.30 per machine-hour Depreciation $68,300 During March, the company worked 18,000 machine-hours and produced 12,000 units. The company had originally planned to work 20,000 machine-hours during March. Required 1. Prepare a flexible budget for March 2. Prepare a report showing the spending variances for March.
Complete this question by entering your answers in the tabs below. Required1Required2 Prepare a flexible budget for March. (Input all amounts as positive values.) FAB Corporation Flexible Budget For the Month Ended March 31 Machine-hours Utilities Maintenance Supplies Indirect labor Depreciation Total
e this question by entering your answers in the tabs below. Required 2 Required 1 Required 2 Prepare a report showing the spending variances for March. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all mounts as positive values) FAB Corporation Spending Variances For the Month Ended March 31 Utilities Maintenance Supplies Indirect labor Depreciation Total
Answer #1

1.

Following shows Preparation of flexible budget for march:

FAB Corporation

Flexible Budget

For the Month Ended March 31

Machine hours 18,000
Utilities ($16,400 + $0.11 \times 18,000) $18,380
Maintenance (38,900 + 2 \times 18,000) 74,900
Supplies (0.5 \times 18,000) 9,000
Indirect labor(94,000 + 1.3 \times 18,000 117,400
Depreciation 68,300
Total $287,980

2.

Spending variance = Flexible budget - Actual results

Following is the Report showing Spending variance for the month of march:

FAB Corporation

Spending Variances

For the Month Ended March 31

Utilities $2,020 "U"
Maintenance 2,200 "F"
Supplies 800 "U"
Indirect labor 3,500 "U"
Depreciation 1,700 "U"
Total $5,820 "U"

> #1 is the Flexible Budget - the Planning Budget to get the activity variances.

Kim Possibley Tue, Apr 19, 2022 7:47 PM